# Bias Detection via Maximum Subgroup Discrepancy

### v1

@@ 5.1 Datasests
To showcase the intended use of our method, we use the US Census data via the Folktables library [12]. Specifically, we use two families of five datasets.
@@
The first family consists of the five pre-defined prediction tasks: ACSIncome, ACSPublicCoverage, ACSMobility, ACSEmployment, and ACSTravelTime. We compare the distributions of samples with $y_{i}=1$ and $y_{i}=0$ on the data of residents of California. For example, in ACSIncome, this means comparing the distributions between people earning more than \$50,000 a year and people earning less, similar to the well-known Adult dataset. For more details see the original paper [12]. To evaluate the distance comparably between methods, we only consider protected attributes, the list of which is in Appendix A.1.1.
@@
The second family of datasets aims at comparing distributions of population between different US states. We take attributes that could be considered protected (e.g., sex, race, disability, age) and take data for five distinct pairs of states. The state pairs are as follows: Hawaii v. Maine, California v. Wyoming, Mississippi v. New Hampshire, Maryland v. Mississippi, and Louisiana v. Utah. The state selection aims to compare demographically, socioeconomically, and culturally different states, allowing for a comparison of diverse populations across the United States. The complete list of used attributes is in Appendix A.1.1.
@@
All used data is 1-year horizon data from the survey year 2018. We work with binary data, where we binarize continuous values to 10 equally spaced bins and semantically reduce the cardinalities of certain subgroups. The sample sizes span from around 10,000 to 380,000 and the number of protected features spans from 3 (with around 1,200 possible subgroups) to 14 (with more than $5\cdot 10^{9}$ subgroups). Details are in Appendix A.1.
5.2 Setup
@@
We aim to see how the number of samples influences the value of the distance. Ideally, one would need only a few samples to compute the true distance of the distributions. For that, each dataset is subsampled to form 5 (smaller) datasets, with sizes forming a geometric sequence from 1000 samples to the maximum number of samples for the given dataset. Each method is evaluated on each dataset. We ran each experiment configuration five times with different random seeds.
Each experiment was run on a cluster. Each node had 32 GB of RAM and either AMD EPYC 7543 or Intel Xeon Scalable Gold 6146 processor. Each distance computation had a time limit of 10 minutes. All code is openly available on GitHub [^3] https://github.com/Epanemu/DNF_bias.
5.3 Results
@@
Since most evaluated methods have different notions of distance, it is difficult to compare them directly. Thus, we present relative distances w.r.t. the best estimate of the distance. In other words, we divide each of the values by the mean distance computed on the entire dataset. Or, if that value cannot be computed, we take the mean value on the second-largest number of samples. For original distance measure comparison, see Figure 3 in Appendix A.2.1.
@@
Figure 2: Mean relative distance measured by each method. The datasets in the top row are standard datasets provided by the Folktables library. The bottom row shows datasets from the US Census, where we consider many protected attributes and compare their distributions of a pair of US states. Dashed lines represent methods that find a subgroup and for which we report the $\mathrm{MSD}_{\mathrm{diff}}$ measure. The color bands represent the standard deviation. Ideally, each relative distance estimate should be 1, meaning that the estimate is the same as the estimate on the highest number of samples.
@@
The means over 5 random seeds are presented in Figure 2. We report Ripper and BRCG results on different numbers of samples because these methods require equal sample sizes for both distributions, so we subsample the distribution for which we have more samples. This is especially pronounced in the California X Wyoming comparison due to the difference in population size. The color bands represent the standard deviations. The BRCG and Ripper often vary in the quality of the distance estimate and have a rather high standard deviation. This is due to the methods being influenced by the randomness of the data. However, nonzero deviation on the last sample size (where data is the same for all seeds) suggests an effect of randomness within the algorithm or the ordering of the data. Clearly, TV and Wasserstein-based measures do not converge quickly to the final value. It is possible that the best estimate is still far from the true distance measure. This shows that we cannot be certain of the true value of the distance measure.
@@
Finally, MMD shows comparable performance to our MIO-based $\mathrm{MSD}_{\mathrm{diff}}$, both in deviation and in convergence. While this seems to empirically point to low sample complexity on these datasets, one cannot be certain that this holds always. In comparison, this is the case for our $\mathrm{MSD}_{\mathrm{diff}}$, with high probability.
@@
The MIO-based $\mathrm{MSD}_{\mathrm{diff}}$ almost always finds the global optimum within 10 minutes. The only exceptions were 13 seeded setups on three datasets with both many protected attributes and many samples, which could be expected. Luckily, we show that we do not need a high number of samples to provide a good estimate of $\mathrm{MSD}_{\mathrm{diff}}$. Additionally, providing the solver with an optimal solution on the smaller sample size could be used to prune some branches in the solver and would likely help with runtime, even on bigger instances. In addition, the strong performance of our method means that we are also able to recover the correct subgroup with maximal distance even with exponentially fewer samples than the maximum over the dataset.
6 Conclusion
We pointed out the difficulties of evaluating the distance of probability distributions by current methods, especially in the context of evaluating bias for intersectional subgroups. We proposed a family of distance measures MSDΔ as the maximal distance $\Delta$ over all subgroups.
@@ 6 Conclusion
We further defined a specific $\mathrm{MSD}_{\mathrm{diff}}$ and proved its linear sample complexity, in contrast to the exponential sample complexities of existing methods. We validated this on datasets of real US Census data with varying numbers of protected attributes and sample sizes. Only MMD showed empirically comparable performance to our MIO-based $\mathrm{MSD}_{\mathrm{diff}}$, but it does not provide as good of an interpretation, that could be useful in bias evaluation.
@@
The MIO formulation also outperforms more general DNF learners, which also lack guarantees due to not being global optimizers. In addition, the global optimality of MIO does not necessarily come at a disadvantage as is otherwise common. That is because the linear sample complexity means that we do not need exponentially many samples, thus the formulation remains smaller and solves faster.
Future work
As we have shown, evaluating bias in subgroups is an intrinsically difficult task. There are many challenges that require more work, one of which is the training of ML models under intersectional fairness constraints. We believe $\mathrm{MSD}_{\mathrm{diff}}$ could be a good measure for learning with fairness guarantees. Additionally, there may be other $\Delta$ functions that could be efficiently used for MSDΔ measures, e.g., the SPSF [27].
Impact Statement
By its nature, the positive potential societal impact of Algorithmic Fairness is well understood, and consequently, the field has been intensively developed in the past decade. The choice of a bias measure has a substantial impact on the evaluation of any AI system. Our bias measure facilitates the study of intersectional fairness of [8], which would be intractable with many other means of evaluation of bias.
@@ Acknowledgment
This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 10107056.
Appendix A Appendix
A.1 Datasets description
We utilize 2 types of datasets, both constructed from the US Census data - American Community Survey (ACS). The first 5 (with names starting with ACS) represent tasks constructed by the creators of the folktables library [12]. We simply access it, taking 1-year survey data of Californians for the year 2018.
The datasets of the other type are constructed by us using the folktables API from the same data source. We select five pairs of distinctive US states and compare their data distributions. We simply take a random half of the available data for each state and combine it to form our dataset. Each dataset has the same attributes, listed in Table 3.
The dataset statistics are in Table 2. Note the Hawaii X Maine dataset for an illustration of the difficulty of assessing the subgroup bias. It is quite difficult to evaluate the bias of more than 400,000x more groups than there are samples.
Table 2: Datasets, sorted by sample size. Datasets starting with ACS are the original tasks, provided by the folktables library [12]. (CA) means that data is taken from the state of California. The bottom half of the datasets are constructed by us from the US Census data, through the folktables library. We compare states using all of the protected attributes from Table 3. Datasets are in the order in which they appear in the plots. The number of subgroups may change even for the same attributes when some categorical value is missing from the data for one dataset/task and not for the other.
Name Number of samples Protected attributes Number of subgroups
@@ A.1 Datasets description
A.1.1 Protected attributes
Below, we list the enumerate the attributes for each dataset. For more discussion and statistics of the attributes, see Table 3.
@@ A.1.1 Protected attributes
Table 3: Protected attributes. Most attributes are commonly considered protected, with multiple being related to race and nationality. We consider Military service protected, due to existing laws against discrimination based on veteran status. The rationale behind protecting the feature of recently giving birth is linked to not uncommon discrimination in the workplace based on the prospect of women being in fertile age. For deeper insight, refer to the official documentation of the ACS data source.
@@
POBP Place of birth 221 6 (grouped by continent)
@@
POVPIP Income / Poverty threshold continuous 10 (binned)
A.2 More results
A.2.1 Original distances
@@
In addition to the results in the main body of the paper, we present the results on the true distances, before “normalization” by the best-estimated distance, in Figure 3. It is indeed difficult to compare the methods together, one might just comment on the distances separately. For example, note the range of estimates of TV, oftentimes, the value changes by 0.3 from the first to last estimate, which is essentially a third of the range of feasible values TV can take.
@@
Also, one can notice that Wasserstein metrics and the MMD are not well-suitable for the evaluation of intersectional bias by comparing results on Hawaii X Maine with Mississippi X New Hampshire. Wasserstein metrics, as well as MMD, evaluate the distance between 0.15 and 0.2 for both datasets, while the most discrepant subgroup has the $\mathrm{MSD}_{\mathrm{diff}}$ distance of cca 0.7 in one and cca 0.3 in the other, less than a half.
Lastly, notice that our proposed MIO-based measure always finds the highest $\mathrm{MSD}_{\mathrm{diff}}$ out of the three optimizers of conjunctions. This is due to the global optimality of the solutions. Out of the other two methods, Ripper seems to be performing better.
Figure 3: Original distances for all methods.
A.2.2 RSE

### v2

@@ 5.1. Datasests
To showcase the intended use of our method, we use the US Census data via the folktables library (Ding et al., 2021). Specifically, we use two families of five datasets.
@@
The first family consists of the five pre-defined prediction tasks provided by the folktables library (Ding et al., 2021). We compare the distributions of samples with $y_{i}=1$ and $y_{i}=0$ on the data of Californians in 2018. For example, in ACSIncome, this means comparing the distributions between people earning more than \$50,000 a year and people earning less, similar to the well-known Adult dataset. For more details, see the original paper (Ding et al., 2021). To make the distance evaluation comparable between methods, we only consider protected attributes. Those are listed in Appendix A.1.1.
@@
The second family of datasets aims to compare population distributions between different US states. We take 14 attributes that could be considered protected (e.g., sex, race, disability, age) and take data for five distinct pairs of states. The state selection aims to compare demographically, socioeconomically, and culturally different states, allowing for a comparison of diverse populations across the United States.
@@
The sample sizes span from around 14,000 to 380,000, and the number of protected features spans from 4 (with around 1,200 possible subgroups) to 14 (with more than $5\cdot 10^{9}$ subgroups). See Table 2 in Appendix A.1 for more details.
5.2. Setup
To solve the proposed MIO formulation (15), we model it using the Pyomo library (Bynum et al., 2021), enabling easy use of a variety of solvers. We use the Gurobi solver (Gurobi Optimization, LLC, 2023). We set the $N_{\text{min}}$ parameter to 10, allowing only subgroups with at least ten samples.
We also implement TV and MMD. In the MMD evaluation, we utilize the overlap kernel function $k(x_{i_{1}},x_{i_{2}})=\sum_{j=1}^{d}\left\llbracket x_{i_{1},j}=x_{i_{2},j}\right\rrbracket/d$ because the considered data is binary or categorical. For W1 and W2, we utilize POT (Flamary et al., 2021), an open-source Python library for the computation of optimal transport.
Since there is no public implementation of the original BRCG, we take the non-MIO re-implementation of BRCG, called BRCG-light (Arya et al., 2019). We use the Ripper algorithm available in the AIX360 library (Arya et al., 2019). We modify both implementations to return a single-term DNF (i.e., a conjunction) to ensure comparability to our method. As addressed earlier, we run Ripper and BRCG two times, once to minimize error and the second time to maximize it (minimize the error with flipped labels). If one of the distributions has more samples, we subsample it to have an equal number of samples for each distribution. This makes minimizing the 0-1 loss equivalent to finding the $\mathrm{MSD}$.
@@ 5.2. Setup
Each dataset is subsampled to form 5 (smaller) datasets, with sizes forming a geometric sequence from 1000 samples to the maximum number of samples for the given dataset. Each method is evaluated on each subsampled dataset. We ran each experiment configuration five times with different random seeds. Each distance computation had a time limit of 10 minutes. All code for replicating the results is openly available on GitHub [^2] https://github.com/Epanemu/MSD. MSD is also implemented as a part of humancompatible.detect package [^3] https://github.com/humancompatible/detect.
5.3. Results
We evaluate three criteria. Firstly, we validate that our approach using MIO improves on naive enumeration of all subgroups (Q1). Secondly, we show that our approach indeed has advantageous sample complexity compared to other distribution distance measures (Q2). Finally, we validate that formulating the $\mathrm{MSD}$ within MIO outperforms standard algorithms for finding DNFs when utilized to calculate $\mathrm{MSD}$ (Q3) and showcase the interpretability of the proposed measure.
5.3.1. Evaluating MS(D)D by enumeration (Q1)
Naively, one could evaluate MSD (or MSDD, recall from Section 1.1) by evaluating some distance for each of the exponential number of subgroups. We performed such an experiment to empirically compare this naive approach to the exponential runtime of solving the formulation (15). We compare to a naive enumeration algorithm to compute $\mathrm{MSD}$ and a few MSDD methods. They systematically enumerate subgroups and compute the given distance function $\Delta$ for each of them, seeking the maximum. We skip all encountered subgroups with fewer than 10 samples to ensure comparability. Subgroups with samples from only one distribution are also skipped when the distance cannot be computed. The enumeration was run on the same hardware, on a single thread in Python. The proposed method is capable of considering exponentially more subgroups in the evaluation of MSD, compared to the naive approach, when given the same time limit; see Figure 2. The $\mathrm{MSD}$ points are on the diagonal since we know we have the maximum over all subgroups when the formulation is solved to global optimality. We omit configurations that were not solved to optimality.
Figure 2. We show the number of subgroups in real datasets and the mean number of subgroups (subsets of the datasets) for which each distance was computed within 10 minutes. The diamonds represent measures computed naively by enumeration, and the squares represent the proposed method.
A log-log plot showing the number of subgroups on the x-axis and the number of considered subgroups within a 10-minute run time on the y-axis. A subgroup is considered if the distance on it was computed or if it has fewer than 10 samples. The proposed method is on the diagonal and dominates other methods that do not reach above 10 to the 7th power threshold.
For the datasets with the most subgroups, even the fastest method manages to consider around 3,000 times fewer subgroups. Extrapolating that, one would need more than 20 days to finish the computation. That is computed on the smallest dataset (Hawaii X Maine), where more than 89% of the “considered” subgroups were skipped due to having too few samples.
5.3.2. Evaluating sample complexity (Q2)
@@
Since most evaluated methods have different notions of distance, it is difficult to compare them directly. Thus, we present relative distances with respect to the best estimate of the distance. In other words, we divide each of the values by the mean distance computed on the entire dataset. Or, if that value cannot be computed, we take the mean value from the second-largest number of samples. For original distance measure comparison, see Figure 5 in Appendix B.1.
@@
Figure 3. Mean relative distance measured by each method. The datasets in the top row are standard datasets provided by the folktables library. The bottom row shows datasets from the US Census, where we consider many protected attributes and compare their distributions on a pair of US states. Dashed lines represent methods that find a subgroup and for which we report the $\mathrm{MSD}$ measure. The color bands represent the standard deviation. Ideally, each relative distance estimate should be 1, meaning that the estimate is the same as the estimate on the highest number of samples.
The figure contains 10 plots in 2 rows; on top are the convergence results for folktables datasets, and on the bottom are results for 5 state pair comparisons. The MIO-based MSD and MMD show good and fast convergence, BRCG and Ripper seem somewhat erratic, and Wasserstein metrics and Total Variation converge, but notably slower.
@@
The means over 5 random seeds are presented in Figure 3. We report Ripper and BRCG results on different numbers of samples because these methods require equal sample sizes for both distributions, so we subsample the distribution for which we have more samples. This is especially pronounced in the California X Wyoming comparison due to the difference in population size. The BRCG and Ripper often vary in the quality of the distance estimate and have a rather high standard deviation. This is due to the methods being influenced by the randomness of the data or the algorithm. Clearly, TV and Wasserstein-based measures do not converge quickly to the final value. It is possible that the best estimate is still far from the true distance measure. This suggests that we cannot be certain of the true value of the distance measure.
@@
Finally, MMD shows comparable performance to our MIO-based $\mathrm{MSD}$, both in deviation and in convergence. While this seems to empirically point to low sample complexity on these datasets, one cannot be certain that this holds always. In contrast, this is the case for our $\mathrm{MSD}$, with high probability.
@@
The MIO-based $\mathrm{MSD}$ almost always finds the global optimum within 10 minutes. The only exceptions were 13 (out of 250) seeded setups on three datasets with many protected attributes and samples, which could be expected. Luckily, we show that we do not need a high number of samples to provide a good estimate of $\mathrm{MSD}$. Additionally, providing the solver with an optimal solution on the smaller sample size could be used to prune some branches in the solver and would likely help with runtime, even on bigger instances. In addition, our method’s strong performance means that we also recover the correct subgroup with maximal distance, even with exponentially fewer samples than the maximum over the dataset.
5.3.3. Comparing algorithms for $\mathrm{MSD}$ (Q3)
As explained in Section 3.2, one can utilize standard DNF (or conjunction) learning algorithms to evaluate $\mathrm{MSD}$. We compare our proposed MIO formulation to Ripper and BRCG specifically, in Figure 4. We run the proposed method on the data with a balanced number of samples (by subsampling) of the two distributions to ensure comparability. The proposed method does not finish within 10 minutes for only three configurations. In all other cases, it returns global optima.
Figure 4. Evaluation of $\mathrm{MSD}$ using three methods. The lines represent mean values, and the color bands represent the standard deviation. MIO-based solutions (ours) dominate the solutions of non-globally optimal solvers. This is not the case for some runs with low sample sizes due to the fact that BRCG and Ripper consider subgroups with fewer than 10 samples.
The figure contains 10 plots in 2 rows; on top are the distance results for folktables datasets, and on the bottom are results for 5 state pair comparisons. The MIO MSD is mostly on top; BRCG and Ripper reach the same value in some plots, but in other plots, they differ by a lot.
The MIO-based $\mathrm{MSD}$ value is always higher than (or equal to) the competitor methods due to the global optimality of the solutions. This does not always hold in some lower-sample configurations, which is due to the fact that BRCG and Ripper were not constrained to consider only subgroups with fewer than 10 samples. More importantly, the $\mathrm{MSD}$ values found by BRCG and Ripper are less stable with increasing sample size and often do not converge to the globally maximal value. This illustrates the unsuitability of BRCG and Ripper, since one does not obtain guarantees.
Furthermore, this guarantee of global maximality enables two things. Firstly, if the MSD is below some given threshold, we can be sure that no subgroup violates the threshold, using the empirical distributions. This threshold can be computed by setting a probability level $\delta$ and utilizing Theorem 3.1 to obtain a guarantee with probability $\delta$. Secondly, we obtain the most disadvantaged subgroup. For example, the group with the highest $\mathrm{MSD}$ for the task of estimating whether the commute to work takes more than 20 minutes (ACSTravelTime) was white women. All found groups are listed in Appendix B.3.
6. Conclusion
@@
We pointed out the difficulties of evaluating the distance of probability distributions by current methods, especially in the context of evaluating bias for intersectional subgroups. We defined the Maximum Subgroup Discrepancy ($\mathrm{MSD}$) and proved its linear sample complexity, in contrast to the exponential sample complexities of existing methods. We validated this on prediction tasks from real US Census data with varying numbers of protected attributes and sample sizes. Only MMD shows empirically comparable sample complexity to our MIO-based $\mathrm{MSD}$, but it does not provide as good an interpretation that could be useful in bias evaluation.
@@
The MIO formulation also outperforms more general DNF learners, which lack guarantees due to not being global optimizers. In addition, the global optimality of MIO does not necessarily come at a disadvantage, as is otherwise common. That is because the linear sample complexity means that we do not need exponentially many samples, thus the formulation remains smaller and solves faster.
@@ Acknowledgment
This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101070568. Access to the computational infrastructure of the OP VVV-funded project CZ.02.1.01/0.0/0.0/16 019/0000765 “Research Center for Informatics” is also gratefully acknowledged. The first author also received funding from the Grant Agency of the Czech Technical University in Prague, grant No. SGS23/184/OHK3/3T/13.
Appendix
Appendix A Experiment details
Each experiment was run on an internal cluster. Each node had 32 GB of RAM and either an AMD EPYC 7543 or an Intel Xeon Scalable Gold 6146 processor, based on availability.
A.1. Datasets description
The dataset statistics are in Table 2. Note the Hawaii X Maine dataset for an illustration of the difficulty of assessing the subgroup bias. It is quite difficult to evaluate the bias of more than 400,000 times more groups than there are samples. We utilize 2 types of datasets, both constructed from the US Census data - American Community Survey (ACS). The first 5 (with names starting with ACS) represent tasks constructed by the creators of the folktables library (Ding et al., 2021). The datasets of the other type are constructed by us using the folktables API from the same data source. We select five pairs of distinctive US states and compare their data distributions. We simply combine a randomly sampled half of the available data for each state to form our dataset. Each dataset has the same attributes listed in Table 3.
All used data is 1-year horizon data from the survey year 2018. We work with categorical or binary data, where we binarize continuous values into 10 equally spaced bins. We semantically reduce the cardinality of the Place of birth attribute from 221 to 6, based on continents (and keeping the US separate).
Table 2. Datasets used in experiments. Those starting with ACS are the original tasks provided by the folktables library (Ding et al., 2021). (CA) means that the data is taken from the state of California. The bottom half of the datasets compares the population of US states using the US Census data. We compare states using 14 protected attributes, listed in Table 3. Datasets are in the order in which they appear in the results. The number of subgroups may change even for the same set of attributes when some categorical value is missing from the data for one dataset/task and not for the other.
Dataset name # samples # protected # subgroups
@@ A.1. Datasets description
A.1.1. Protected attributes
Below, we enumerate the attributes for each dataset. More details about the attributes are in Table 3.
@@ A.1.1. Protected attributes
Table 3. Protected attributes. Most attributes are commonly considered protected, with multiple being related to race and nationality. We consider Military service protected, due to existing laws against discrimination based on veteran status. The rationale behind protecting the feature of recently giving birth is linked to not uncommon discrimination in the workplace based on the prospect of women being of fertile age. For deeper insight, refer to the official documentation of the ACS data source.
@@
POBP Place of birth 221 6 (grouped to continents and the USA)
@@
POVPIP Income, relative to the Poverty threshold continuous 10 (binned)
Appendix B Further results
Figure 5. Original distances for all methods.
The figure contains 10 plots in 2 rows; on top are the convergence results for folktables datasets, and on the bottom are results for 5 state pair comparisons. The proposed MSD and MMD are the only ones with stable performance.
B.1. Original distances
@@
In addition to the results in the main body of the paper, we present the results on the true distances, before “normalization” by the best-estimated distance, in Figure 5. It is difficult to compare the methods due to the various distances having different interpretations. One might just comment on the distances separately. For example, note the range of estimates of TV; oftentimes, the value changes by around 0.3 from the first to the last estimate, which is essentially a third of the range of feasible values TV can take.
@@
Also, one can notice that Wasserstein metrics and the MMD are not well-suited for the evaluation of maximal intersectional bias. When comparing results on Hawaii X Maine with Mississippi X New Hampshire. Wasserstein metrics and MMD estimates differ by around 0.05 between the two datasets, while $\mathrm{MSD}$ distance differs by around 0.4, which is more than half of the value for Hawaii X Maine. In other words, $\mathrm{MSD}$ reports vastly different distances for the two datasets, one being twice as close, while Wasserstein metrics and MMD report the two datasets having much more similar distances, differing at most by a quarter, in the case of MMD. This is expected since $\mathrm{MSD}$ finds a maximally (dis)advantaged subgroup.
Finally, notice the difference in distance estimates between the ACSIncome and the Louisiana X Utah setups. The $\mathrm{MSD}$ estimate is higher in the second case, while Wasserstein metrics, much like MMD, report higher distance in the first case. In practice, utilizing the Wasserstein or MMD could lead to choosing a classifier that seems less biased, but has a subgroup with notably higher disparity compared to all subgroup disparities of the rejected classifier.
Figure 6. Relative standard error of all methods.
The figure contains 10 plots in 2 rows; on top are the relative standard errors for folktables datasets, and on the bottom are errors for 5 state pair comparisons. The Wasserstein metrics and Total Variation have the lowest errors throughout, BRCG and Ripper have errors somewhat unstable, and the MSD and MMD have similar errors, decreasing with the number of samples.
B.2. RSE